OpenStudy (anonymous):

plesae it is an emergency I need help with the problem andrew can paint the neighbor's house 2 times as fast as bailey. The year Andrew and bailey worked together, it took 1 days. How long would it take each to paint the house? I need help on this Please

jimthompson5910 (jim_thompson5910):

1/x + 1/(2x) = 1/t 1/x + 1/(2x) = 1/1 1/x + 1/(2x) = 1 2/(2x) + 1/(2x) = 1 (2+1)/(2x) = 1 3/(2x) = 1 3 = 1*2x 3 = 2x 3/2 = x x = 3/2 x = 1 1/2 So it will take Baily 1 and half days to paint the house and it will take Andrew 2*(3/2) = 3 days to paint the house

OpenStudy (anonymous):

thankssssssssssssssssssssssssssss